# NeRF-based Point Cloud Reconstruction using a Stationary Camera for Agricultural Applications

### v1

@@ Abstract
This paper presents a NeRF-based framework for point cloud (PCD) reconstruction, specifically designed for indoor high-throughput plant phenotyping facilities. Traditional NeRF-based reconstruction methods require cameras to move around stationary objects, but this approach is impractical for high-throughput environments where objects are rapidly imaged while moving on conveyors or rotating pedestals. To address this limitation, we develop a variant of NeRF-based PCD reconstruction that uses a single stationary camera to capture images as the object rotates on a pedestal. Our workflow comprises COLMAP-based pose estimation, a straightforward pose transformation to simulate camera movement, and subsequent standard NeRF training. A defined Region of Interest (ROI) excludes irrelevant scene data, enabling the generation of high-resolution point clouds (10M points). Experimental results demonstrate excellent reconstruction fidelity, with precision-recall analyses yielding an F-score close to 100.00 across all evaluated plant objects. Although pose estimation remains computationally intensive with a stationary camera setup, overall training and reconstruction times are competitive, validating the method’s feasibility for practical high-throughput indoor phenotyping applications. Our findings indicate that high-quality NeRF-based 3D reconstructions are achievable using a stationary camera, eliminating the need for complex camera motion or costly imaging equipment. This approach is especially beneficial when employing expensive and delicate instruments, such as hyperspectral cameras, for 3D plant phenotyping. Future work will focus on optimizing pose estimation techniques and further streamlining the methodology to facilitate seamless integration into automated, high-throughput 3D phenotyping pipelines.
@@
Figure 1: Schematic of the stationary camera imaging system for NeRF-based point cloud reconstruction in high-throughput plant phenotyping. In this setup, each plant is conveyed to a rotating turntable marked against a matte black background. Over a full 30-second rotation, a tripod-mounted stationary camera captures high-resolution images that serve as input for NeRF techniques to generate 3D reconstructions. This streamlined approach eliminates the need for complex moving-camera rigs, aligning with the objectives of efficient, scalable agricultural imaging. The right shows different PCD reconstruction using the stationary camera. (a) Apricot, (b) Banana, (c) Bell pepper, (d) Maize ear, (e) Crassula ovata, and (f) Haworthia sp..
@@ 1 Introduction
Conventional approaches to 3D phenotyping primarily involve photogrammetry techniques such as structure-from-motion (SfM) and multi-view stereo (MVS)[8, 5], as well as terrestrial laser scanning (TLS)[23]. Although these methods provide detailed structural data and have been effectively applied across various crops [16, 19], they present several practical limitations, including high equipment costs, manual labor, and significant computational demands [1, 29]. Additionally, their scalability and capacity to capture minute structural details in dynamic agricultural scenarios are limited [25, 22].
@@
To address these limitations, this paper presents a stationary-camera-based NeRF framework explicitly developed for indoor high-throughput phenotyping. Our approach integrates COLMAP-based pose estimation and a simple pose transformation to simulate camera movement, enabling standard NeRF training using stationary-camera data. We demonstrate the effectiveness and scalability of our method, achieving high-fidelity point cloud reconstructions with near-perfect precision-recall metrics. Our primary contributions include: (1) a novel stationary-camera NeRF reconstruction pipeline designed specifically for high-throughput indoor phenotyping, (2) extensive experimental validation demonstrating reconstruction fidelity, and (3) evidence of computational feasibility, paving the way for seamless integration into automated phenotyping workflows.
@@ 2 Background
Building on these advances, our work proposes a three-channel approach tailored for agricultural applications. Our goal is to develop a method that is robust, low-cost, and high-throughput by optimizing image resolution, reducing the number of required images, and ensuring accurate color representation. This approach minimizes setup constraints while delivering high-fidelity 3D reconstructions, making it ideally suited for real-world agricultural scenarios.
@@ 3.1 Experimental Setup
For pose estimation and structured scene reconstruction, 5×5 ArUco markers were generated using the ArUco Markers Generator (https://chev.me/arucogen/). The markers were positioned in two configurations: six markers were attached to a 3D-printed blue cylinder (diameter = 0.09 m, height = 0.07 m), which was designed using Onshape (a cloud SaaS CAD solution) and fabricated with an Ultimaker S5 3D printer. The cylinder was designed to fit into a sphere in CloudCompare for scaling, ensuring accurate dimension matching in the reconstruction process. Additionally, eight markers were placed on a circular paper mounted on the turntable for pose estimation and structured scene reconstruction. These markers provided feature points for COLMAP, enabling accurate camera extrinsic computation.
A ping pong ball (radius = 0.04 m) was included in the scene for metric scale calibration, measured using a digital caliper to ensure accurate dimension scaling.
@@ 4 Results and Discussion
The experimental results demonstrate that the proposed stationary camera setup can produce high-quality PCD reconstructions, as illustrated in Figure 1(right). The reconstructed models closely align with the ground-truth PCD obtained using the rotating camera setup. The primary evaluation metrics included precision-recall analysis (Figure 4) and computation time comparisons for pose estimation, training, and reconstruction (Table 1).
@@
The PCD reconstruction times between SC and GT were comparable, with differences of only a few minutes across all datasets. The largest discrepancy occurred in Crassula ovata (5m57.6s for SC vs. 5m3.5s for GT), showing a 54.1s difference, while banana exhibited the smallest difference (3m54.9s for SC vs. 4m57.8s for GT). These results suggest that although pose estimation is computationally intensive in SC, the overall workflow remains feasible and competitive in terms of training and reconstruction efficiency.
@@ 5 Conclusions
This paper introduces a NeRF-based point cloud (PCD) reconstruction framework explicitly designed for indoor high-throughput plant phenotyping environments. Traditional NeRF methods require cameras to move around stationary objects, a process incompatible with automated indoor phenotyping facilities that routinely employ stationary cameras alongside rotating pedestals or conveyors. To overcome this limitation, we develop a stationary-camera-based NeRF approach that simulates camera motion through a straightforward pose transformation after COLMAP-based pose estimation, facilitating standard NeRF training.
Our experimental validation demonstrated that this method achieves high-fidelity 3D reconstructions, yielding precision-recall F-scores approaching 100% across various agriculturally relevant objects. Despite the computational time associated with pose estimation in stationary setups, our framework showed competitive overall reconstruction times, highlighting its practical feasibility for integration into automated phenotyping workflows.
Key contributions of our work include: (1) introducing a stationary-camera NeRF-based pipeline tailored explicitly for high-throughput indoor phenotyping, (2) providing experimental evidence supporting reconstruction fidelity and consistency, and (3) establishing the computational practicality of this approach, even when using high-resolution image data.
This work has direct relevance for plant science research, breeding programs, and agricultural production, particularly when employing expensive or fragile imaging instruments such as hyperspectral cameras. By eliminating the need for complex camera rigs and costly imaging hardware, our method simplifies high-throughput phenotyping infrastructure, reducing both operational complexity and expenses.
Future research directions will involve optimizing pose estimation processes to further reduce computational demands, exploring the method’s adaptability to a broader range of plant species and phenotyping scenarios, and integrating multimodal imaging data such as RGB and hyperspectral imaging.

### v2

# SC-NeRF: NeRF-based Point Cloud Reconstruction using a Stationary Camera for Agricultural Applications

@@ Abstract
This paper presents a NeRF-based framework for point cloud (PCD) reconstruction, specifically designed for indoor high-throughput plant phenotyping facilities. Traditional NeRF-based reconstruction methods require cameras to move around stationary objects, but this approach is impractical for high-throughput environments where objects are rapidly imaged while moving on conveyors or rotating pedestals. To address this limitation, we develop a variant of NeRF-based PCD reconstruction that uses a single stationary camera to capture images as the object rotates on a pedestal. Our workflow comprises COLMAP-based pose estimation, a straightforward pose transformation to simulate camera movement, and subsequent standard NeRF training. A defined Region of Interest (ROI) excludes irrelevant scene data, enabling the generation of high-resolution point clouds (10M points). Experimental results demonstrate excellent reconstruction fidelity, with precision-recall analyses yielding an F-score close to 100.00 across all evaluated plant objects. Although pose estimation remains computationally intensive with a stationary camera setup, overall training and reconstruction times are competitive, validating the method’s feasibility for practical high-throughput indoor phenotyping applications. Our findings indicate that high-quality NeRF-based 3D reconstructions are achievable using a stationary camera, eliminating the need for complex camera motion or costly imaging equipment. This approach is especially beneficial when employing expensive and delicate instruments, such as hyperspectral cameras, for 3D plant phenotyping. Future work will focus on optimizing pose estimation techniques and further streamlining the methodology to facilitate seamless integration into automated, high-throughput 3D phenotyping pipelines. We provide all datasets and our code, available at https://baskargroup.github.io/SC-NeRF/
@@
Figure 1: Schematic of the stationary camera imaging system for NeRF-based point cloud reconstruction in high-throughput plant phenotyping. In this setup, each plant is conveyed to a rotating turntable marked against a matte black background. Over a full 30-second rotation, a tripod-mounted stationary camera captures high-resolution images that serve as input for NeRF techniques to generate 3D reconstructions. This streamlined approach eliminates the need for complex moving-camera rigs, aligning with the objectives of efficient, scalable agricultural imaging. The right shows different PCD reconstruction using the stationary camera. (a) Apricot, (b) Banana, (c) Bell pepper, (d) Maize ear, (e) Crassula ovata, and (f) Haworthia sp.
@@ 1 Introduction
Conventional approaches to 3D phenotyping primarily involve photogrammetry techniques such as structure-from-motion (SfM) and multi-view stereo (MVS)[8, 5], as well as terrestrial laser scanning (TLS)[23, 33]. Although these methods provide detailed structural data and have been effectively applied across various crops [16, 19], they present several practical limitations, including high equipment costs, manual labor, and significant computational demands [1, 29]. Additionally, their scalability and capacity to capture minute structural details in dynamic agricultural scenarios are limited [25, 22].
@@
To address these limitations, this paper presents a stationary-camera-based NeRF framework explicitly developed for indoor high-throughput phenotyping. Unlike moving-camera setups, our approach uses a rotating pedestal, cutting cost and complexity while supporting sensitive tools like hyperspectral cameras. Our approach integrates COLMAP-based pose estimation and a simple pose transformation to simulate camera movement, enabling standard NeRF training using stationary-camera data. We demonstrate the effectiveness and scalability of our method, achieving high-fidelity point cloud reconstructions with near-perfect precision-recall metrics. Our primary contributions include: (1) a novel stationary-camera NeRF reconstruction pipeline designed specifically for high-throughput indoor phenotyping, (2) extensive experimental validation demonstrating reconstruction fidelity, and (3) evidence of computational feasibility, paving the way for seamless integration into automated phenotyping workflows.
@@ 2 Background
Building on these advances, our work proposes a three-channel approach tailored for agricultural applications. Our goal is to develop a method that is robust, low-cost, and high-throughput by optimizing image resolution, reducing the number of required images, and ensuring accurate color representation. This approach minimizes setup constraints, such as those required by OB-NeRF [32], while delivering high-fidelity 3D reconstructions, making it ideally suited for real-world agricultural scenarios.
@@ 3.1 Experimental Setup
For pose estimation and structured scene reconstruction, 5×5 ArUco markers were generated using the ArUco Markers Generator (https://chev.me/arucogen/). The markers were positioned in two configurations: six markers were attached to a 3D-printed blue cylinder (diameter = 0.09 m, height = 0.07 m), which was digitally designed and 3D-printed using an FDM-based desktop printer. A custom cylindrical container was designed to enhance visual clarity and facilitate reliable marker tracking during reconstruction. The container featured a uniformly colored surface to minimize artifacts caused by blending with the black background. Its constant radius ensured smooth and continuous visibility of the attached markers, addressing the geometric inconsistencies of standard plant pots. In addition, eight markers were placed on a circular paper mounted to the turntable to maximize pose estimation support. This was particularly important for the stationary camera setup, where only the object and markers change between frames—unlike rotating camera environments, which provide natural parallax and viewpoint diversity. These markers provided consistent feature points for COLMAP, enabling accurate computation of camera extrinsic parameters and contributing to robust structured scene reconstruction. A ping pong ball (radius = 0.04 m) was included in the scene for metric scale calibration, measured using a digital caliper to ensure accurate dimension scaling.
@@ 4 Results and Discussion
The experimental results demonstrate that the proposed stationary camera setup can produce high-quality PCD reconstructions, as illustrated in Figure 1(right). The reconstructed models closely align with the ground-truth PCD obtained using the rotating camera setup. The primary evaluation metrics included precision-recall analysis (Figure 4) and computation time comparisons for pose estimation, training, and reconstruction (Table 1). The dataset created for this work can be found at https://huggingface.co/datasets/BGLab/SC-NeRF.
@@
The PCD reconstruction times between SC and GT were comparable, with differences of only a few minutes across all datasets. The largest discrepancy occurred in Banana (4m57.8s for GT vs. 3m54.9s for SC), showing a 63.8s difference. In contrast, the smallest difference occurred in Apricot (5m4.4s for GT vs. 4m32.0s for SC), with a gap of only 32.4s. These results suggest that although pose estimation is computationally intensive in SC, the overall workflow remains feasible and competitive in terms of training and reconstruction efficiency.
@@
Dataset Availability:
To support reproducibility, we release the full dataset used in this study as SC-NeRF. It contains all intermediate and final outputs from our NeRF-based reconstruction pipeline for six agriculturally relevant objects: apricot, banana, bell pepper, maize ear, Crassula ovata, and Haworthia sp.. Each object was recorded under two settings: SC: Stationary camera, rotating object (our method). GT: Moving camera, stationary object (used as ground truth). The dataset is organized as follows:
raw/: Includes 4K videos (.MOV) and extracted frames.
pre/: Contains COLMAP camera poses and sparse reconstructions, formatted via ns-process-data.
train/: Holds trained NeRF model checkpoints using the nerfacto trainer.
pcd/: Final aligned and scaled 10M-point clouds, filtered via outlier removal.
Each file is named according to the object and capture mode (e.g., banana_sc, apricot_gt). The dataset can support training, benchmarking, or extending NeRF-based 3D reconstruction pipelines in agriculture. The dataset can be found at https://huggingface.co/datasets/BGLab/SC-NeRF.
@@ 5 Conclusions
This paper introduces a NeRF-based point cloud (PCD) reconstruction framework explicitly designed for indoor high-throughput plant phenotyping environments. Traditional NeRF methods require cameras to move around stationary objects, a process incompatible with automated indoor phenotyping facilities that routinely employ stationary cameras alongside rotating pedestals or conveyors. To overcome this limitation, we develop a stationary-camera-based NeRF approach that simulates camera motion through a straightforward pose transformation after COLMAP-based pose estimation, facilitating standard NeRF training. Our experimental validation demonstrated that this method achieves high-fidelity 3D reconstructions, yielding very high precision-recall F-scores across various agriculturally relevant objects. Despite the computational time associated with pose estimation in stationary setups, our framework showed competitive overall reconstruction times, highlighting its practical feasibility for integration into automated phenotyping workflows.
This work has direct relevance for plant science research, breeding programs, and agricultural production, particularly when employing expensive or fragile imaging instruments such as hyperspectral cameras. By eliminating the need for complex camera rigs and costly imaging hardware, our method simplifies high-throughput phenotyping infrastructure, reducing both operational complexity and expenses. Future research directions will involve optimizing pose estimation processes to reduce computational demand, exploring the method’s adaptability to a broader range of plant species and phenotyping scenarios, and integrating multimodal data such as hyperspectral bands with RGB imaging.